OpenStudy (anonymous):

Simplify the expression. (Do not expand the expression in the denominator.) ((e^(7x) + e^(-7x) (e^(7x) + e^(-7x) - (e^(7x) - e^(-7x)(e^(7x) - e^(-7x)) / ((e^(7x) + e^(-7x))^2

OpenStudy (anonymous):

Can you type it in latex? You are missing a few parenthesis :P

OpenStudy (anonymous):

latex?

OpenStudy (anonymous):

(e^(7x) + e^(-7x))^2 - (e^(7x) - e^(-7x))^2 is the numerator ((e^(7x) + e^(-7x))^2 is the denominator

OpenStudy (anonymous):

Okay so you have: \[\frac{(e^{7x}+e^{-7x})^2-(e^{7x}-e^{-7x})^2}{(e^{7x}+e^{-7x})^2}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay. Start off my multiplying the top out completely. \[(e^{7x}+e^{-7x})^2=e^{7x}*e^{7x}+e^{-7x}*e^{-7x}+e^{-7x}*e^{7x}+e^{-7x}*e^{7x}=e^{14x}+e^{-14x}+2\]

OpenStudy (anonymous):

The last one should end as: \[e^{14x}+e^{-14x}+2\] Then the second part is the exact same except with a negative sign. so: \[(e^{7x}-e^{-7x})^2=-e^{7x}*e^{-7x}+e^{7x}*e^{7x}+e^{-7x}*e^{-7x}-e^{7x}*e^{-7x}\] Which equals: \[e^{14x}+e^{-14x}-2\]

OpenStudy (anonymous):

Distribute the negative on the second term giving: \[e^{14x}+e^{-14x}+2-e^{14x}-e^{-14x}+2=4\]

OpenStudy (anonymous):

So now you have: \[\frac{4}{(e^{7x}+e^{-7x})^2}\]

OpenStudy (anonymous):

Ahhhh coolies, thanks a heap!!!

OpenStudy (anonymous):

If the writing looks bad you can refresh the page.

OpenStudy (anonymous):

Nope, looks good XD

OpenStudy (anonymous):

No problem :P Sorry it took a bit to type up. Latex is the way I was writing the math btw :P It just looks nice instead of all the text parenthesis and what not.

OpenStudy (anonymous):

Yeah it does :)